# Combinatorial Selection with Costly Information

### v1

@@ 3 An Amortization Framework
In this section, we develop a novel cost amortization technique that will allow us to lower bound the cost of the optimal adaptive algorithm for matroid-min-CICS. We first establish our amortization for the special case of Markov chains (Section 3.1). We then connect our amortization framework to the notion of optimality curves (Section 3.2) and use this connection to extend our approach to general MDPs (Section 3.3). Omitted proofs are presented in Appendix A.
@@ 3.1 Amortized Surrogate Costs for Markov Chains
Let $\mathbb{I}=(\mathcal{M}_{1},\cdots,\mathcal{M}_{n},\mathcal{F})$ be an instance of matroid-min-CICS where every MDP $\mathcal{M}_{i}$ is a Markov chain. In this setting, an algorithm chooses an index $i\in[n]$ at every step and advances $\mathcal{M}_{i}$ along the unique action available. It is well known that the optimal algorithm is an index policy: we associate every state $s_{i}\in S_{i}$ for $i\in[n]$ with an index, and advance the chain with the minimum index at every step. Importantly, the index is independent of the evolution of states in the algorithm, and depends solely on the description of the Markov chain. The policy can therefore be computed easily. We will now describe a new, simple proof of the optimality of index-based policies based on an amortization of action costs.
We begin with some notation. For a Markov chain $\mathcal{M}=(S,\sigma,A,c,\mathcal{D},V,T)$ and a state $s\in S$, we use $c(s)$ to denote the cost of the unique action in $A(s)$. We use $T(s)\subseteq T$ to denote the set of terminal states reachable from $s$. Consider an algorithm that starts in $\sigma$ and advances the chain $\mathcal{M}$ along the unique available action until it reaches a terminal state. For $s\in S$, let $p(s)$ denote the probability that the state $s$ is encountered during this process. Observe that by definition $p(s)=\sum_{t\in T(s)}p(t)$; $p(\sigma)=1$; and $p$ defines a probability distribution over the terminal states in $T$. We are now ready to define our notion of cost amortization.
Definition 3.
A cost amortization of a Markov chain $\mathcal{M}=(S,\sigma,A,c,\mathcal{D},V,T)$ is a non-negative vector $b=\{b_{st}\}_{s\in S,t\in T(s)}$ with the property that $\sum_{t\in T(s)}p(t)b_{st}=p(s)c(s)$ for all states $s\in S$. Based on this amortization, we define:
  • The amortized cost of a terminal state $t\in T$ as $\rho_{b}(t):=v(t)+\sum_{s:t\in T(s)}b_{st}$.
  • The surrogate cost of the Markov chain $\mathcal{M}$ as the random variable $\rho_{\mathcal{M},b}$ that takes on value $\rho_{b}(t)$ for $t\in T$ with probability $p(t)$.
  • The index of a state $s\in S$ of the Markov chain $\mathcal{M}$ as $I_{\mathcal{M},b}(s)=\min_{t\in T(s)}\rho_{b}(t)$.
Observe that only the terminal states carry amortized costs; essentially, the amortization distributes the cost of every action across its downstream terminals. The surrogate cost of a Markov chain is the amortized cost of the random terminal state realized by running the chain until it terminates.
Borrowing terminology from Singla (2017), we can relate the performance of any algorithm on $\mathbb{I}$ to its performance in a “free information world” where the algorithm does not pay action costs but, whenever it accepts at a terminal state, is responsible for paying the entire amortized cost at that state. Since the algorithm does not accept every Markov chain it advances, we immediately get the following lower bound.
@@
Consider a matroid-min-CICS $\mathbb{I}=(\mathcal{M}_{1},\cdots,\mathcal{M}_{n},\mathcal{F})$ over Markov chains and let $b_{i}$ be any cost amortization of $\mathcal{M}_{i}$ with surrogate cost $\rho_{i}:=\rho_{\mathcal{M}_{i},b_{i}}$ for all $i\in[n]$. Then, the expected cost of any algorithm for $\mathbb{I}$ is at least $\operatorname{E}\left[\min_{S\in\mathcal{F}}\sum_{i\in S}\rho_{i}\right]$.
We will now exhibit a specific cost amortization and a corresponding algorithm that achieves the lower bound of Lemma 3.1, proving optimality. The water filling cost amortization is described algorithmically in a bottom-up fashion. Recall that each state in the Markov chain is associated with a unique sequence of actions that lead to it; let the length of this sequence denote the “level” of the state. We start from the terminal states, and define their total cost share to be equal to their value $v(t)$. We then proceed over non-terminal states in decreasing order of level. Each non-terminal state $s\in S$ distributes its total cost $c(s)$ across its downstream terminals $T(s)$, starting from the terminals with the lowest current total cost, until the equation $\sum_{t\in T(s)}p(t)b_{st}=p(s)c(s)$ is satisfied. We use $W^{*}_{\mathcal{M}}$ to denote the water filling surrogate cost of a Markov chain $\mathcal{M}$, and $I^{*}_{\mathcal{M}}(s)$ to denote the water filling index of a state $s$ in $\mathcal{M}$.
@@
Consider an MDP $\mathcal{M}$ whose starting state is given a choice between following one of two Markov chains $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$. The Markov chain $\mathcal{M}_{1}$ has a single action of cost $c_{1}=1$, leading to a terminal of value $v(t_{11})=(2/3)$ with probability $p_{11}=(3/4)$ and to a terminal of value $v(t_{12})=4$ with probability $p_{12}=(1/4)$. The Markov chain $\mathcal{M}_{2}$ has a single action of cost $c_{2}=(1/8)$, leading to a terminal of value $v(t_{21})=(1/2)$ with probability $p_{12}=(1/4)$ and to a terminal of value $v(t_{22})=3$ with probability $p_{22}=(3/4)$. We pictorially represent $\mathcal{M}$ as well as the water filling amortizations for $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ in Figure 1.
Figure 1: The water filling amortization for Markov chains $\mathcal{M}_{1}$ (left) and $\mathcal{M}_{2}$ (right). For each chain, the water filling amortization computes a unique index $g$ (here, $g_{1}=2$ and $g_{2}=1$), corresponding to the value for which the highlighted area equals the cost of the amortized action. For each terminal $t$, the corresponding cost share is given by $(g-v(t))^{+}$; here, $b_{11}=4/3$, $b_{21}=1/2$ and $b_{12}=b_{22}=0$. Therefore, the surrogate cost $W^{*}_{\mathcal{M}_{1}}$ is $2$ with probability $(3/4)$ and $4$ with probability $(1/4)$. Likewise, $W^{*}_{\mathcal{M}_{2}}$ is $1$ with probability $(1/4)$ and $3$ with probability $(3/4)$.
We now describe an index-based policy for matroid-min-CICS based on the water filling amortization:
Definition 4.
The Water Filling Index policy for an instance $\mathbb{I}=(\mathcal{M}_{1},\cdots,\mathcal{M}_{n},\mathcal{F})$ of matroid-min-CICS chooses at every step the Markov chain $i^{*}=\operatorname{argmin}_{i\in\mathcal{F}_{S}}I^{*}_{i}(s_{i})$, where $s_{i}$ is the current state of Markov chain $\mathcal{M}_{i}$; $S$ is the set of terminated (selected) Markov chains; and $\mathcal{F}_{S}=\{i:\text{rank}(S\cup\{i\})>\text{rank}(S)\}$.
The water filling amortization ensures that if $b_{st}>0$ for some state $s\in S$ and terminal $t\in T(s)$, then the index of all states in the unique path from $s$ to $t$ will be the same. Since the water filling index policy always advances the Markov chain of minimum index, this implies that for any action taken by the policy, any downstream terminal that “owes” a non-zero cost share to that action will be accepted if instantiated. Therefore all the amortized cost shares are paid in expectation, establishing the following optimality result.
Theorem 3.2.
For any matroid-min-CICS instance $\mathbb{I}=(\mathcal{M}_{1},\cdots,\mathcal{M}_{n},\mathcal{F})$ over Markov chains, the expected cost of the water filling index policy is equal to $\operatorname{E}\left[\min_{S\in\mathcal{F}}\sum_{i\in S}W^{*}_{\mathcal{M}_{i}}\right]$. The policy is therefore optimal for instance $\mathbb{I}$.
@@ 3.2 Local Games and Optimality Curves
We now connect the water filling amortization described above to the notion of optimality curves for MDPs defined in the work of Whittle (1980) and its follow ups. We first define a “local game” for an MDP $\mathcal{M}$.
Definition 5.
The local game $(\mathcal{M},y)$ is a single-selection min-CICS with two MDPs, one of which is the MDP $\mathcal{M}$. The second MDP, a.k.a. the outside option, is a deterministic option with a single state of value $y$ and a single “accept” action available. A policy for the local game advances $\mathcal{M}$ for some (zero or non-zero) number of steps and either accepts the deterministic option $y$ or the value from $\mathcal{M}$. Let $f_{\mathcal{M}}(y)$ denote the expected cost of the optimal policy for the local game $(\mathcal{M},y)$. We refer to the function $f_{\mathcal{M}}$ as the Optimality Curve of the MDP $\mathcal{M}$.
@@
Definition 6.
Let $\mathcal{M}$ be an MDP with optimality curve $f_{\mathcal{M}}$. The Water Filling Surrogate Cost for $\mathcal{M}$ is the random variable $W^{*}_{\mathcal{M}}$ generated by picking a value from the CDF $1-\frac{d}{dy}f_{\mathcal{M}}(y)$. That is, $W^{*}_{\mathcal{M}}$ is the random variable satisfying $f_{\mathcal{M}}(y)=\operatorname{E}\left[\min(y,W^{*}_{\mathcal{M}})\right]$ for all $y\in\real$.
Unfortunately, this definition does not give us insight into how the water filling surrogate costs relate to the structure of the MDP, or whether and how we can extend Lemma 3.1 or Theorem 3.2 to arbitrary MDPs. We will now address these issues by directly relating optimality curves to the water filling procedure.
@@ 3.3 Water Filling and Surrogate Costs for General MDPs
In the previous sections, we established that in the case of Markov chains, the water filling surrogate cost of a randomly sampled terminal state recovers the optimality curve. The challenge to extending this argument for general MDPs is that each sequence of actions creates a different distribution over the terminal states. We address this with the following lemma:
@@
For any MDP $\mathcal{M}=(S,\sigma,A,c,\mathcal{D},V,T)$, there exists an amortized cost function $\rho:T\mapsto\Delta(\real)$ mapping terminal states to distributions over costs and a non-negative cost sharing vector $b=\{b_{st}\}_{s\in S,t\in T(s)}$, such that for all committing policies $\pi\in\mathcal{C}\left(\mathcal{M}\right)$, generating a Markov chain $\mathcal{M}^{\pi}$ with states $S_{\pi}\subseteq S$, terminal states $T_{\pi}\subseteq T\cap S_{\pi}$, and a distribution $p_{\pi}$ over them, the following properties hold:
@@
    Action Independence. The water filling surrogate cost $W^{*}_{\mathcal{M}}$ corresponds to sampling a terminal state $t\sim p_{\pi}$ and then sampling from distribution $\rho(t)$.
@@
    Cost Sharing. For all $t\in T_{\pi}$, it holds that $\operatorname{E}\left[\rho(t)\right]=v(t)+\sum_{s\in S_{\pi}:t\in T_{\pi}(s)}b_{st}$.
@@
    Cost Dominance. For all $s\in S_{\pi}$, it holds that $\sum_{t\in T_{\pi}(s)}p_{\pi}(t)b_{st}\leq p_{\pi}(s)c(a)$ for the unique action $a\in A(s)$ chosen by $\pi$.
Lemma 3.3 allows us to express the water filling surrogate cost of the MDP through a cost sharing and amortization much in the same way as Definition 3 does for Markov chains. Each terminal is assigned a random amortized cost that is independent of the trajectory of the algorithm. Yet, as the action independence condition states, every committing policy generates the same distribution over amortized costs, matching the water filling surrogate costs defined based on optimality curves. This allows us to account for the eventual cost of any algorithm without worrying about the specific actions it takes.
The lemma also defines action independent cost shares paid by terminals to the upstream actions. However, the action independence comes at a cost. In contrast to the cost shares defined for Markov chains in Definition 3, these cost shares only recover a part of the cost of the action. As a result, water filling surrogate costs provide us with a lower bound, but not an exact accounting, of the cost of any algorithm for the CICS. The proof of the following lower bound closely mirrors the lower bound we previously established for Markov chains (Lemma 3.1).
Theorem 3.4.
For any matroid-min-CICS instance $\mathbb{I}=(\mathcal{M}_{1},\cdots,\mathcal{M}_{n},\mathcal{F})$, the expected cost of the optimal adaptive policy is at least $\operatorname{E}\left[\min_{S\in\mathcal{F}}\sum_{i\in S}W^{*}_{\mathcal{M}_{i}}\right]$.
Figure 2: We consider the same setting as in Example 1. The optimality curve for MDP $\mathcal{M}$ is obtained by the minimum of the optimality curves for $\mathcal{M}_{1}$ (orange line) and $\mathcal{M}_{2}$ (blue line). From the optimality curve, we see that $W^{*}_{\mathcal{M}}$ is $1$ with probability $(1/4)$, $2.5$ with probability $(1/2)$ and $4$ with probability $(1/4)$. From Lemma 3.3, $W^{*}_{\mathcal{M}}$ can be obtained by first sampling from either $W^{*}_{\mathcal{M}_{1}}$ or $W^{*}_{\mathcal{M}_{2}}$ and then applying a randomized mapping $\rho_{\mathcal{M}}(\cdot)$ to the sampled value. The mapping probabilities are shown in the table.
We conclude this section with the proof of Lemma 3.3.
@@
Fix any MDP $\mathcal{M}=(S,\sigma,A,c,\mathcal{D},V,T)$. The proof follows by induction on the horizon $H$ of the MDP. If $H=1$, then the starting state $\sigma$ is the unique terminal state of $\mathcal{M}$ and thus all three conditions are satisfied by the trivial mapping $\rho(\sigma):=v(\sigma)$. We now assume that the conditions of the lemma hold for all MDPs of horizon up to $H$ and extend it to MDPs of horizon $H+1$.
Let $\mathcal{M}$ be of horizon $H+1$ and let $A(\sigma)=\{a_{1},\cdots,a_{k}\}$ be the set of actions available at the starting state $\sigma$. Let $R_{j}$ denote the set of all states that can be reached directly by taking action $a_{j}$ on $\sigma$; that is, $R_{j}:=\{s\in S:\mathcal{D}(\sigma,a_{j},s)>0\}$. For each $s\in R_{j}$, we use $\mathcal{M}_{s}$ to denote the subprocesses of $\mathcal{M}$ that starts in state $s$ and $W^{*}_{\mathcal{M}_{s}}$ to denote the water filling surrogate cost of $\mathcal{M}_{s}$. By definition, each MDP $\mathcal{M}_{s}$ has horizon up to $H$ and thus by the induction hypothesis it admits a mapping $\rho_{s}(\cdot)$ over its terminal states $T(s)$ and a non-negative cost sharing vector $b^{s}$ satisfying the properties of the lemma.
We will define the surrogate cost function $\rho$ in two steps. First we will compute the water filling amortization of each action $a_{j}$. Let $Z_{j}$ denote the random variable that first draws a state $s$ from the distribution $\mathcal{D}(\sigma,a_{j},\cdot)$ and then draws a value from the distribution $W^{*}_{\mathcal{M}_{s}}$. Let $g_{j}$ be the solution to the equation
@@
$=\min\left\{y,\min_{j\in[k]}\bigg{(}c(a_{j})+\operatorname{E}\left[\min\{y,Z_{j}\}\right]\bigg{)}\right\}$
$=\min\left\{y,\min_{j\in[k]}\operatorname{E}\left[\min\bigg{(}y,\hat{Z}_{j}\bigg{)}\right]\right\}.$

### v2

@@ 3 An Amortization Framework
In this section, we develop a novel cost amortization technique that will allow us to lower bound the cost of the optimal adaptive algorithm for matroid-min-CICS. We first establish our amortization for the special case of Markov chains (Section 3.1). We then connect our amortization framework to the notion of optimality curves (Section 3.2) and use this connection to extend our approach to general MDPs (Section 3.3). All proofs are deferred to Appendix D, except for the proof of Lemma 3.5 which can be found at the end of this section.
@@ 3.1 Amortized Surrogate Costs for Markov Chains
In this section we consider instances $\mathbb{I}=(\mathbb{M},\mathcal{F})$ of the matroid-min-CICS where every MDP $\mathcal{M}_{i}$ is a Markov chain. We begin with some notation. For a Markov chain $\mathcal{M}=(S,\sigma,A,c,\mathcal{D},V,T)$ and a state $s\in S$, we use $c(s)$ to denote the cost of the unique action in $A(s)$. A trajectory $\tau$ in this Markov chain is a sequence of states that results from advancing the chain until it terminates. For a state $s\in S$, let $\Upsilon(s)$ denote the set of all possible trajectories starting in $s$ and let $\Upsilon:=\Upsilon(\sigma)$. For state $s\in S$ and trajectory $\tau\in\Upsilon(s)$, let $p(\tau)$ denote the probability that a random trajectory starting in the initial state $\sigma$ visits the state $s$ and then continues along $\tau$ until terminating. Let $p(s):=\sum_{\tau\in\Upsilon(s)}p(\tau)$ denote the probability of visiting $s$. For $\tau\in\Upsilon$ and $s\in\tau$, let $\tau_{s}$ denote the suffix of $\tau$ starting with $s$. Finally, let $v(\tau)$ be the value $v(t)$ of the terminal state $t$ that $\tau$ ends in.
We will now define an amortization of action costs in a Markov chain over the resulting (random) trajectories that follow. The intention is that in the amortized setting, the algorithm does not incur any action costs when the action is taken but pays the amortized cost of the instantiated trajectory only if it follows the trajectory all the way to termination.
Definition 4 (Cost Amortization).
A cost amortization of a Markov chain $\mathcal{M}=(S,\sigma,A,c,\mathcal{D},V,T)$ is a non-negative vector $b=\{b_{s\tau}\}_{s\in S,\tau\in\Upsilon(s)}$ with the property that $\sum_{\tau\in\Upsilon(s)}p(\tau)b_{s\tau}=p(s)c(s)$ for all states $s\in S$. Based on this amortization, we define:
  • The amortized cost of a trajectory $\tau\in\Upsilon$ as $\rho_{b}(\tau):=v(\tau)+\sum_{s\in\tau}b_{s\tau_{s}}$.
  • The surrogate cost of the Markov chain $\mathcal{M}$ as the random variable $\rho_{\mathcal{M},b}$ that takes on value $\rho_{b}(\tau)$ for $\tau\in\Upsilon$ with probability $p(\tau)$.
  • The index of a state $s\in S$ of the Markov chain $\mathcal{M}$ as $I_{\mathcal{M},b}(s):=\min_{\tau\in\Upsilon:s\in\tau}\rho_{b}(\tau)$.
Note that the “cost shares” $b_{s\tau}$ distributed by a state $s$ towards its resulting trajectories fully cover the cost of $s$’s action and no more. However, in the amortized setting, these cost shares are paid only when the trajectory terminates, and go unpaid if the algorithm stops advancing the Markov chain midway through the trajectory. We therefore obtain the following lemma.
@@
Consider any instance $\mathbb{I}=(\mathbb{M},\mathcal{F})$ of matroid-min-CICS over Markov chains and let $b_{i}$ be any cost amortization of $\mathcal{M}_{i}$ with surrogate cost $\rho_{i}:=\rho_{\mathcal{M}_{i},b_{i}}$ for all $i\in[n]$. Then, $\mathrm{OPT}(\mathbb{I})\geq\operatorname{E}\left[\min_{S\in\mathcal{F}}\sum_{i\in S}\rho_{i}\right]$.
We will now exhibit a specific cost amortization and a corresponding algorithm that achieves the lower bound of Lemma 3.1, proving optimality. The water filling cost amortization is described algorithmically in a bottom-up fashion. We consider the states of the chain in reverse topological order, starting from the terminals up towards the root $\sigma$. Trajectories $\tau\in\Upsilon(t)$ for terminal states $t\in T$ are singletons and do not carry cost shares. Consider a state $s$ such that the cost shares for all states reachable from it have been determined. The state distributes its total cost $c(s)$ across its downstream trajectories $\tau\in\Upsilon(s)$, starting from those with the lowest current total cost, until the equation $\sum_{\tau\in\Upsilon(s)}p(\tau)b_{s\tau}=p(s)c(s)$ is satisfied. Formally, let $\tau^{\prime}=\tau\setminus s$ be the sequence of states in $\tau$ following $s$; its current downstream cost is $\rho(\tau^{\prime})=v(\tau^{\prime})+\sum_{s^{\prime}\in\tau^{\prime}}b_{s^{\prime}\tau^{\prime}_{s^{\prime}}}$. We find the lowest water level $g$ such that the cost shares $b_{s\tau}:=(g-\rho(\tau\setminus s))^{+}$ satisfy the cost equation above. The new total cost of $\tau$ becomes $\rho(\tau)=b_{s\tau}+\rho(\tau\setminus s)=\max\{g,\rho(\tau\setminus s)\}$. We continue in this manner until $\sigma$’s cost is amortized.
We use $W^{*}_{\mathcal{M}}$ to denote the water filling surrogate cost of a Markov chain $\mathcal{M}$, and $I^{*}_{\mathcal{M}}(s)$ to denote the water filling index of a state $s$ in $\mathcal{M}$. Finally, we can describe an index-based policy based on the water filling amortization.
Definition 5 (Water Filling Index Policy).
The water filling index policy for an instance $\mathbb{I}=(\mathbb{M},\mathcal{F})$ of matroid-min-CICS over Markov chains selects at every step the Markov chain $i^{*}=\operatorname{argmin}_{i\in\mathcal{F}_{S}}I^{*}_{i}(s_{i})$, where $s_{i}$ is the current state of Markov chain $\mathcal{M}_{i}$; $S$ is the set of terminated (selected) Markov chains; and $\mathcal{F}_{S}=\{i:\text{rank}(S\cup\{i\})>\text{rank}(S)\}$.
Note that, by definition, the indices of states encountered on any trajectory weakly increase with each next action. The water filling amortization ensures that if $b_{s\tau}>0$ for some state $s\in S$ and $\tau\in\Upsilon(s)$, then the index of all states in the downstream trajectory $\tau$ stays equal to the index of $s$. Since the policy described above always advances the Markov chain of minimum index, this implies that for any action taken by the policy, any downstream trajectory that “owes” a non-zero cost share to that action will terminate with acceptance if instantiated. In other words, all the amortized cost shares are paid in expectation, establishing the following optimality result:
Theorem 3.2.
For any matroid-min-CICS instance $\mathbb{I}=(\mathbb{M},\mathcal{F})$ over Markov chains, the expected cost of the water filling index policy is equal to $\operatorname{E}\left[\min_{S\in\mathcal{F}}\sum_{i\in S}W^{*}_{\mathcal{M}_{i}}\right]$. The policy is therefore optimal for $\mathbb{I}$.
To conclude this section, we will illustrate water filling through an example.
@@
Consider an MDP $\mathcal{M}$ whose starting state is given a choice between following one of two Markov chains $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$. Each Markov chain has a single costly action and results in a distribution over two distinct states, as illustrated in Figure 2. The action costs, terminal values, and instantiation probabilities are indicated in the figure. For each chain, the water filling amortization computes a unique index $g$ (here, $g_{1}=2$ and $g_{2}=1$), corresponding to the value for which the highlighted area equals the cost of the amortized action. Each trajectory leads to a unique terminal node, so we can associate cost shares with terminals. For each terminal $t$, the corresponding cost share is given by $(g-v(t))^{+}$; here, $b_{11}=4/3$, $b_{21}=1/2$ and $b_{12}=b_{22}=0$. Therefore, the surrogate cost $W^{*}_{\mathcal{M}_{1}}$ is $2$ with probability $(3/4)$ and $4$ with probability $(1/4)$. Likewise, $W^{*}_{\mathcal{M}_{2}}$ is $1$ with probability $(1/4)$ and $3$ with probability $(3/4)$.
Figure 2: The water filling amortization for Markov chains $\mathcal{M}_{1}$ (left branch) and $\mathcal{M}_{2}$ (right branch). The solid black edges denote costly actions (of costs $1$ and $1/8$) and the dashed blue edges denote random transitions to terminal states upon taking these actions, with the corresponding probabilities of transition. Each terminal state has a corresponding value, depicted in the bottom line. The graphs on the right display the water-filling operation for each chain with the $x$-axis representing probabilities of instantiation of the respective trajectories and the $y$-axis representing the terminal values for these trajectories. The area of the gray region equals the action cost and its height is the water level that determines cost shares. By definition of water-filling, cost shares are first transferred to the trajectories of minimum value.
@@ 3.2 Local Games and Optimality Curves
We now connect the water filling amortization described above to the notion of optimality curves for MDPs, defined in the work of Whittle (1980) and its follow ups. We first define a “local game” for an MDP $\mathcal{M}$.
Definition 6 (Local Game and Optimality Curve).
The local game $(\mathcal{M},y)$ is a single-selection min-CICS with two MDPs, one of which is the MDP $\mathcal{M}$. The second MDP, a.k.a. the outside option, has a single terminal state with deterministic value $y$. A policy for the local game advances $\mathcal{M}$ for some (zero or non-zero) number of steps and either accepts the deterministic option $y$ or the value from $\mathcal{M}$. Let $f_{\mathcal{M}}(y)$ denote the expected cost of the optimal policy for the local game $(\mathcal{M},y)$. We refer to the function $f_{\mathcal{M}}$ as the optimality curve of the MDP $\mathcal{M}$.
@@
Definition 7 (Water Filling Surrogate Costs).
Let $\mathcal{M}$ be an MDP with optimality curve $f_{\mathcal{M}}$. The water filling surrogate cost for $\mathcal{M}$ is the random variable $W^{*}_{\mathcal{M}}$ generated from the CDF $1-\frac{d}{dy}f_{\mathcal{M}}(y)$. That is, $W^{*}_{\mathcal{M}}$ is the random variable satisfying $f_{\mathcal{M}}(y)=\operatorname{E}\left[\min(y,W^{*}_{\mathcal{M}})\right]$ for all $y\in\real$.
Our eventual goal is to prove the following analog of Theorem 3.2 for general CICS over MDPs.
Theorem 3.4.
In any instance $\mathbb{I}=(\mathbb{M},\mathcal{F})$ of matroid-min-CICS, $\mathrm{OPT}(\mathbb{I})\geq\operatorname{E}\left[\min_{S\in\mathcal{F}}\sum_{i\in S}W^{*}_{\mathcal{M}_{i}}\right]$.
The quantity $\operatorname{E}\left[\min_{i}W^{*}_{\mathcal{M}_{i}}\right]$ is called the Whittle integral in the context of single selection over infinite-horizon discounted-reward MDPs. It was first suggested as a lower bound on the optimal cost by Whittle (1980) and proved in that context by Brown and Smith (2013) and Hadfield-Menell and Russell (2015). Unfortunately, Definition 7 does not give us insight into how the surrogate costs $W^{*}_{\mathcal{M}_{i}}$ relate to the cost of an adaptive algorithm for matroid-min-CICS. We will instead prove the theorem by relating optimality curves to the water filling procedure.
@@ 3.3 Water Filling and Surrogate Costs for General MDPs
In this section we will prove Theorem 3.4. In the case of Markov chains, the water filling surrogate cost of a randomly sampled trajectory recovers the optimality curve of the Markov chain. The challenge to extending this argument for general MDPs is that each sequence of actions creates a different distribution over the trajectories and consequently also over the surrogate costs, as illustrated in Example 1.
Our main observation is that if we are willing to cover action costs only partially through the amortization, then we can define an appropriate amortization for every possible deterministic commitment in $\mathcal{M}$ such that the distribution over surrogate costs generated by that commitment is independent of the commitment itself. In other words, every sequence of actions in the MDP generates the same distribution over surrogate costs. This distribution is a fundamental property of the MDP itself, and not of the algorithm operating on the MDP. We formalize this property through the following lemma.
@@
For any MDP $\mathcal{M}=(S,\sigma,A,c,\mathcal{D},V,T)$ and any deterministic commitment $\pi\in\mathcal{C}\left(\mathcal{M}\right)$, generating a Markov chain $\mathcal{M}^{\pi}$ with states $S_{\pi}\subseteq S$, realizable trajectories $\Upsilon_{\pi}\subseteq\Upsilon$ and a distribution $p_{\pi}$ over trajectories and reachable states, there exists an amortized cost function $\rho^{\pi}:\Upsilon_{\pi}\mapsto\Delta(\real)$ mapping trajectories to distributions over costs and a non-negative cost sharing vector $b^{\pi}=\{b^{\pi}_{s\tau}\}_{s\in S_{\pi},\tau\in\Upsilon_{\pi}(s)}$, such that the following properties hold:
@@
    Cost Sharing: the amortized cost of a trajectory pays for its own acceptance value and the cost shares it sends to upstream states. For all $\tau\in\Upsilon_{\pi}$, it holds that $\operatorname{E}\left[\rho^{\pi}(\tau)\right]=v(\tau)+\sum_{s\in\tau}b^{\pi}_{s\tau_{s}}$.
@@
    Cost Dominance: the cost shares received by any state pay towards its action cost, but do not overpay. For all $s\in S_{\pi}$, it holds that $\sum_{\tau\in\Upsilon_{\pi}(s)}p_{\pi}(\tau)b^{\pi}_{s\tau}\leq p_{\pi}(s)c(\pi(s))$.
@@
    Action Independence: sampling a trajectory $\tau\sim p_{\pi}$ and then sampling from the amortized cost distribution $\rho^{\pi}(\tau)$ generates a random surrogate cost for the MDP. This random variable is distributed identically to the water filling surrogate cost $W^{*}_{\mathcal{M}}$.
As mentioned earlier, the amortized costs $\rho^{\pi}(\tau)$ defined by the lemma are necessarily different from the water filling costs in the respective Markov chains $\mathcal{M}^{\pi}$ as they must satisfy action independence. Observe that in the absence of the last condition connecting the surrogate costs $\rho^{\pi}(\tau)$ to the water filling costs $W^{*}_{\mathcal{M}}$ as defined in Definition 7, the lemma is trivial to satisfy. Indeed we can define all of the cost shares to be 00 and still satisfy cost dominance and action independence. The key contribution of the lemma is to show that we can achieve these properties while recovering the “optimal” surrogate costs as determined by the optimality curve $f_{\mathcal{M}}$. Figure 3 illustrates how the lemma applies to the MDP in Example 1.
Once we establish the lemma, the proof of Theorem 3.4 follows in much the same way as the proof of Lemma 3.1. At a high level, we can view the algorithm’s choices in hindsight as corresponding to some deterministic commitment, and relate the algorithm’s cost to the surrogate costs as instantiated through that commitment; the details are deferred to the appendix. We conclude this section with the proof of Lemma 3.5. A casual reader can safely skip this proof and proceed to Section 4.
Figure 3: We consider the same setting as in Example 1. The optimality curve for MDP $\mathcal{M}$ is obtained by the minimum of the optimality curves for $\mathcal{M}_{1}$ (orange dashed line) and $\mathcal{M}_{2}$ (blue dashed line). Each action/commitment leads to two possible trajectories. The table on the right displays the four trajectories along with their respective probabilities of instantiation, values, and amortized costs under their respective commitments (commitment 1 for the first two columns and commitment 2 for the last two columns). Observe that resulting surrogate costs have different distributions. The last row in the table defines an alternate set of (random) surrogate costs $\rho_{\mathcal{M}}(t)$. Observe that for each $t$, we have $\operatorname{E}[\rho_{\mathcal{M}}(t)]\leq v(t)$. This implies cost dominance. The distribution of $\rho_{\mathcal{M}}(t)$ when $t$ is picked according to commitment 1 is $1$ with probability $(3/4)\cdot(1/3)=(1/4)$, $2.5$ with probability $(3/4)\cdot(2/3)=(1/2)$ and $4$ with probability $(1/4)\cdot(1)=(1/4)$. We obtain the same distribution if $t$ is picked according to commitment 2. The same distribution also arises from the solid black optimality curve in the graph on the left, exhibiting the action independence property.
@@
Fix any MDP $\mathcal{M}=(S,\sigma,A,c,\mathcal{D},V,T)$. The proof follows by induction on the horizon $H$ of the MDP. If $H=1$, then $\mathcal{M}$ is a singleton MDP where $\sigma\in T$ is the unique state, $\pi_{\sigma}$ is the unique (trivial) commitment and $\Upsilon_{\pi_{\sigma}}=\{\{\sigma\}\}$. Therefore, the lemma is immediately satisfied by $\rho^{\pi_{\sigma}}(\{\sigma\}):=v(\sigma)$ and $b^{\pi_{\sigma}}_{\sigma\{\sigma\}}:=0$. We now assume that the conditions of the lemma hold for all MDPs of horizon up to $H$ and extend it to MDPs of horizon $H+1$.
Let $\mathcal{M}$ be of horizon $H+1$ and let $A(\sigma)=\{a_{1},\cdots,a_{k}\}$ be the set of actions available at the starting state $\sigma$. Let $R_{j}$ denote the set of all states that can be reached directly by taking action $a_{j}$ on $\sigma$; that is, $R_{j}:=\{s\in S:\mathcal{D}(\sigma,a_{j},s)>0\}$. For each $s\in R_{j}$, we use $\mathcal{M}_{s}$ to denote the subprocesses of $\mathcal{M}$ that starts in state $s$ and $W^{*}_{\mathcal{M}_{s}}$ to denote the water filling surrogate cost of $\mathcal{M}_{s}$. We begin by computing the water filling amortization of each action $a_{j}$.
Let $Z_{j}$ denote the random variable that first draws a state $s$ from the distribution $\mathcal{D}(\sigma,a_{j},\cdot)$ and then draws a value from the distribution $W^{*}_{\mathcal{M}_{s}}$. Let $g_{j}$ be the solution to the equation
@@
$=\min\left\{y,\min_{j\in[k]}\bigg{(}c(a_{j})+\operatorname{E}\left[\min\{y,Z_{j}\}\right]\bigg{)}\right\}=\min\left\{y,\min_{j\in[k]}\operatorname{E}\left[\min\bigg{(}y,\hat{Z}_{j}\bigg{)}\right]\right\}.$